OpenStudy (some.random.cool.kid):

Two rockets, A and B, are shot from two different launch pads. The path of rocket A can be represented by the quadratic function A(t) = −(t − 10)2 + 455, where height, A(t), is in meters, and time, t, is in seconds. The path of rocket B is shown below.

OpenStudy (some.random.cool.kid):

Which rocket reaches the highest altitude (in meters) and how much higher does it travel than the other rocket? Rocket A reaches the highest altitude and it travels 10 meters higher. Rocket B reaches the highest altitude and it travels 10 meters higher. Rocket A reaches the highest altitude and it travels 70 meters higher. Rocket B reaches the highest altitude and it travels 70 meters higher.

OpenStudy (some.random.cool.kid):

@ParthKohli @sammixboo

OpenStudy (some.random.cool.kid):

@rishavraj

OpenStudy (some.random.cool.kid):

@Owlcoffee

OpenStudy (some.random.cool.kid):

@nincompoop

OpenStudy (some.random.cool.kid):

@Catseyeglint911 @imqwerty @whpalmer4

OpenStudy (some.random.cool.kid):

@phi

OpenStudy (some.random.cool.kid):

@♪Chibiterasu can you help?

OpenStudy (phi):

what is the highest that rocket B reaches, based on the graph ? any idea ?

OpenStudy (some.random.cool.kid):

I believe that the highest is just about 500? give or take... @phi

OpenStudy (some.random.cool.kid):

@paki

OpenStudy (phi):

you can read off the number , because they label the peak. what does it say?

OpenStudy (some.random.cool.kid):

uh- 20,525?

OpenStudy (some.random.cool.kid):

is height b the 525?

OpenStudy (phi):

that is an (x,y) pair the 20 is the x-axis (or time) the y value is the height yes 525

OpenStudy (phi):

for A A(t) = −(t − 10)2 + 455 the peak of a parabola (or quadratic as they call it here) is at its vertex this equation is in vertex form, so you can read off the vertex (if you know how... it should be in your notes)

OpenStudy (some.random.cool.kid):

give me a minute ill tell you if i get it or not.

OpenStudy (some.random.cool.kid):

notes - "The vertex form of a quadratic function is given by f (x) = a(x - h)2 + k, where (h, k) is the vertex of the parabola." @phi ok? so... im not sure, am i going to be dividin the vertex by something or compare it to height A?

OpenStudy (phi):

you match up f (x) = a(x - h)2 + k with your equation A(t) = −(t − 10)2 + 455 just focus on the 10 and 455

OpenStudy (phi):

the idea is 10 matches up with the h and the 455 with the k the vertex is at (h,k) so you put in those numbers and the height is the "y" value of the (h, k) pair. i.e. the 2nd number

OpenStudy (some.random.cool.kid):

Which rocket reaches the highest altitude (in meters) and how much higher does it travel than the other rocket.. what i have so far - {A(t) = −(t − 10)2 + 455}.. What i need 'distance between height' i have the (h) value for height A right? so dont i have to solve something to find height B? [or] is 10 height A and 455 is height B? thats where im confused...

OpenStudy (phi):

in (h,k) the h is the time and k is the height so for A, what is its height ?

OpenStudy (some.random.cool.kid):

I thought it was 20 to be honest, yes?

OpenStudy (some.random.cool.kid):

or 10?

OpenStudy (some.random.cool.kid):

still confused... whats confusing is that im using what i see on the graph..

OpenStudy (phi):

How did you get 20? you match up f (x) = a(x - h)2 + k with your equation A(t) = −(t − 10)2 + 455 The graph is rocket B. We could figure out its equation (but we don't know it off hand) the equation A(t) = −(t − 10)2 + 455 is for rocket A we don't have its graph (but we could plot it) so we have to use two different ideas to get the height: for rocket B we look for the vertex and read off the number (easy to do, they tell us the vertex is at (20,525) and the y value (height) is the 2nd number) for rocket A, we have to use the equation to read off the vertex. so, what is the (x,y) or in other words (h, k) for rocket A using the *equation*?

OpenStudy (some.random.cool.kid):

OOOOHH see, thats where i got the 20 from.

OpenStudy (phi):

so far, from the graph we know rocket B reached 525 now we have to use the equation for A to find its height

OpenStudy (some.random.cool.kid):

they already gave us rocket B? makes sense now.

OpenStudy (phi):

they gave us a graph for B. You still had to figure out that the max height was at 525. they also give us an equation for A, we still have to figure out its vertex (which you can do if you know how)

OpenStudy (some.random.cool.kid):

well, on a honest note perhaps you can show me how, and then I will try and solve it myself using whatever techniques you've got, eh?

OpenStudy (phi):

your notes say ***notes - "The vertex form of a quadratic function is given by f (x) = a(x - h)2 + k, where (h, k) is the vertex of the parabola." **** can you match up f (x) = a(x - h)2 + k A(t) = −(t − 10)2 + 455 what does "h" match up with what does "k" match up with ?

OpenStudy (some.random.cool.kid):

uh h -10? and k 455?

OpenStudy (phi):

almost. notice that the (x-h) matches with (t-10) the x /t match up the - /- match up the h / 10 match in other words h matches with 10 (not -10) that means the vertex (h, k) is at (10,455)

OpenStudy (some.random.cool.kid):

ok

OpenStudy (phi):

now that you know the vertex for rocket A, you know its max height

OpenStudy (some.random.cool.kid):

max is 455?

OpenStudy (some.random.cool.kid):

if so do i subtract A and B's max hieghts?

OpenStudy (phi):

yes, remember in (10, 455) the first number is how far sideways we go, and the 2nd number is how far up or down. we want the up/down number

OpenStudy (some.random.cool.kid):

heights*

OpenStudy (phi):

The very first thing I do is write down what we know A 455 B 525 which is higher ? B! then I would look at the choices and cross off any that say A is higher

OpenStudy (some.random.cool.kid):

so rocket b reaches highest at altitude 70, yes?

OpenStudy (phi):

you mean B goes 70 meters higher than A in other words if we go to 455 meters then go another 70 meters 455+70= 525 meters , we get to B's height.

OpenStudy (some.random.cool.kid):

sorry couldnt paste so i wrote it in my own words... meant to say "Rocket B reaches the highest altitude and it travels 70 meters higher."

OpenStudy (phi):

yes, that is good.

OpenStudy (some.random.cool.kid):

thanks you :D